OpenStudy (firejay5):

Simplify each expression. 4. 3m^3 - 3m over 6m^4 * 4m^5 over m + 1

OpenStudy (mertsj):

\[\frac{3m^3-3m}{6m^4}\times\frac{4m^5}{m+1}\]

OpenStudy (mertsj):

Pull 3m out of the first numerator

OpenStudy (firejay5):

yes

OpenStudy (mertsj):

\[\frac{3m(m^2-1)}{6m^4}\times\frac{4m^5}{m+1}\]

OpenStudy (mertsj):

Is that what you got?

OpenStudy (firejay5):

yes

OpenStudy (mertsj):

Now factor m^2-1

OpenStudy (firejay5):

(m-1) (m+1)

OpenStudy (mertsj):

\[\frac{3m(m-1)(m+1)}{6m^4}\times\frac{4m^5}{m+1}\]

OpenStudy (mertsj):

Now cancel all common factors

OpenStudy (firejay5):

cancelled out the (m + 1)

OpenStudy (mertsj):

What about cancelling 3?

OpenStudy (mertsj):

What about 2?

OpenStudy (firejay5):

what do you mean?

OpenStudy (mertsj):

How about m^4?

OpenStudy (mertsj):